OpenStudy (anonymous):

During a recent trip to the store, a cars velocity went from 0 to 60 mph in 20 seconds. what is the average acceleration of the car in mile per hour per hour?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

surely a lot easier than the last one!

OpenStudy (anonymous):

okay thanks, just had to make sure i was right

OpenStudy (anonymous):

average a = delta v/t (t in hours) =60*3600/20 =10,800 mh^-2